OpenStudy (anonymous):

What is the length of the missing side of the triangle in simplest radical form?

OpenStudy (anonymous):



OpenStudy (anonymous):

@ganeshie8 can you help me please

ganeshie8 (ganeshie8):

it looks like a right triangle, use pythagora theorem

ganeshie8 (ganeshie8):

a^2 + b^2 = c^2

ganeshie8 (ganeshie8):

6^2 + 10^2 = c^2

ganeshie8 (ganeshie8):

solve c

OpenStudy (anonymous):

its 2 radical 34

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

good job !